OpenStudy (anonymous):

Write the equation of the line, in point-slope form, that contains the points (-2, -2) and (2, 2).

OpenStudy (anonymous):

@Bookworm14

OpenStudy (anonymous):

did you find the slope?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

lol yeah

OpenStudy (anonymous):

it is \((-2,-2),(2,2)\) right ?

OpenStudy (anonymous):

Mhm

OpenStudy (anonymous):

from -2 to 2 is right 4, and from -2 to 2 is up four right four, up four slope is \(\frac{4}{4}=0\)

OpenStudy (anonymous):

actually it is \[\frac{4}{4}=1\] not zero

OpenStudy (anonymous):

also notice that in both coordinates the x and the y are the same

OpenStudy (anonymous):

And?

OpenStudy (anonymous):

this tells you the equation is \(y=x\)

OpenStudy (anonymous):

Okay, I still need the entire equation.

OpenStudy (bookworm14):

to find slope you must use the formula \[m = \frac{ y2-y1 }{x2-x1 }\] so for this it would be 2- (-2) / 2- (-2) equaling 4/4 or 1 point-slope is in the form of y-y1 = m(x-x1) so for this just substitute your points y- (-2)= 1 (x-x1) now distribute on the "x" side y+2=1x+2 subtract 2 from both sides y=1x or just y=x

OpenStudy (anonymous):

is there something not entire in \(\huge y=x\)?